### v1

# Capacity-Aware Inference: Mitigating the Straggler Effect in Mixture of Experts

Abstract
The Mixture of Experts (MoE) is an effective architecture for scaling large language models by leveraging sparse expert activation, optimizing the trade-off between performance and efficiency. However, under expert parallelism, MoE suffers from inference inefficiencies due to imbalanced token-to-expert assignment, where some experts are overloaded while others remain underutilized. This imbalance leads to poor resource utilization and increased latency, as the most burdened expert dictates the overall delay, a phenomenon we define as the Straggler Effect. To mitigate this, we propose Capacity-Aware Inference, including two key techniques: (1) Capacity-Aware Token Drop, which discards overloaded tokens to regulate the maximum latency of MoE, and (2) Capacity-Aware Token Reroute, which reallocates overflowed tokens to underutilized experts, balancing the token distribution. These techniques collectively optimize both high-load and low-load expert utilization, leading to a more efficient MoE inference pipeline. Extensive experiments demonstrate the effectiveness of our methods, showing significant improvements in inference efficiency, e.g., 0.2% average performance increase and a 1.94$\times$ inference speedup on Mixtral-8$\times$7B-Instruct.
Capacity-Aware Inference: Mitigating the Straggler Effect in
Mixture of Experts
Shwai He1 Weilin Cai2 Jiayi Huang2 Ang Li1 1University of Maryland, College Park 2The Hong Kong University of Science and Technology (Guangzhou) shwaihe@umd.edu, angliece@umd.edu
1 Introduction
In recent years, the rapid evolution of Large Language Models (LLMs) OpenAI (2024); Team (2024a); DeepSeek-AI et al. (2024b) has driven a wave of innovations, continuously expanding the frontiers of AI research and applications. Among the model architectural innovations, the Mixture of Experts (MoE) framework has emerged as a pivotal technique for optimizing the cost-performance trade-off in LLMs. Specifically, MoE Shazeer et al. (2017a); Fedus et al. (2022) enhances scalability by integrating multiple experts while activating only a subset per input. This selective activation substantially improves model performance without a corresponding increase in computational cost, effectively balancing efficiency and performance.
Despite the success of MoE, a key efficiency challenge lies in the imbalanced token-to-expert distribution, which results in some experts being overloaded while others remain underutilized Lepikhin et al. (2021); Zoph et al. (2022). In the context of expert parallelism, low-load experts complete their computations faster but must wait for high-load experts to finish, leading to inefficient resource utilization and increased latency. As illustrated in Figure 1, this phenomenon is defined as the Straggler Effect, where the most burdened expert dictates the overall latency of the MoE layer.
Figure 1: Illustration of the Straggler Effect in MoE Inference, where the most burdened expert (i.e., Straggler Expert) dictates the overall latency.
While auxiliary balance losses have been incorporated into the training process to alleviate imbalance Shazeer et al. (2017a); Fedus et al. (2022); DeepSeek-AI et al. (2024b), these techniques remain ineffective in mitigating imbalance during inference. Specifically, as shown in Figure 1, our findings reveal a highly uneven token distribution among experts, with the highest-load expert handling more than seven times the expected average load. Moreover, managing such an imbalance during inference often incurs additional resource costs. For instance, DeepSeek-V3 addresses this issue by duplicating high-load experts and deploying them redundantly DeepSeek-AI et al. (2024b). This motivates us to explore efficient token-to-expert assignment through two key questions: (1) How to prevent extreme overloading in high-load experts? and (2) How can we utilize the available capacity of low-load experts?
We address these challenges through Capacity-Aware Inference. For high-load experts, we introduce Capacity-Aware Token Drop, a technique that removes excessive tokens from overloaded experts. On the one hand, eliminating these excess tokens alleviates extreme load imbalances, significantly enhancing efficiency. On the other hand, dropping these tokens minimally impacts model performance, as they constitute only a small fraction of the total tokens. For low-load experts, we extend Token Drop with Capacity-Aware Token Reroute, which leverages the available capacity of underutilized experts to process overflowed tokens. Token Reroute further enhances the performance of Token Drop within the capacity-constrained inference framework. Extensive experimental results validate the effectiveness of our proposed techniques, demonstrating significant improvements in both efficiency and performance, e.g., 0.2% average performance increase and a 1.94$\times$ inference speedup in Mixtral-8$\times$7B-Instruct. In short, our contributions are in three folds:
  • We identify the Straggler Effect caused by token imbalance at inference time in Mixture of Experts, highlighting the optimization potential for reducing latency.
  • Toward token imbalance, we propose Token Drop and Token Reroute to enhance the balanced utilization of experts.
  • Experimental results demonstrate the effectiveness of Token Drop and Token Reroute, achieving significant improvements in expert utilization and inference efficiency.
2 Related Works
Mixture of Experts Models
The Mixture of Experts (MoE) is a kind of neural network architecture with an extended set of parameters (referred to as “experts”) controlled by a router, which is first introduced in the context of conditional computation (Jacobs et al., 1991; Jordan and Jacobs, 1994). The potential of sparse activation in MoE is subsequently exploited by Shazeer et al. (2017b) for efficient training and inference on pretrained models with special designs, opening the door for MoE in various vision (Riquelme et al., 2021) and language (Lepikhin et al., 2020; Du et al., 2022; Fedus et al., 2022) scenarios. Attributed to its exceptional efficiency, MoE has been adopted as a foundational framework in the designs of large language models (Jiang et al., 2024; Dai et al., 2024; Xue et al., 2024a; Zhu et al., 2024; Team, 2024b), achieving superior scaling laws at low computational costs. Despite these advancements, MoE still faces efficiency challenges in both training and inference Cai et al. (2024), and our work specifically focuses on enhancing inference-time efficiency.
Imbalance in Mixture of Experts
The imbalance in token-to-expert assignments Zhou et al. (2022); Chen et al. (2022) poses a significant challenge to the deployment of Mixture of Experts (MoE). This imbalance leads to inefficiencies in memory, computation, and communication (He et al., 2023; Song et al., 2023; Xue et al., 2024b), making it a critical bottleneck for MoE scalability and deployment. To mitigate this issue, an auxiliary balance loss Shazeer et al. (2017a) is incorporated into the training process to encourage more uniform token distribution across experts. Additionally, various training strategies have been introduced to further balance token assignments: Switch-Transformer Fedus et al. (2022) and DeepSeek-V2 DeepSeek-AI et al. (2024a) implement Token Drop to alleviate expert overload, while DeepSeek-V3 DeepSeek-AI et al. (2024b) introduces an additional sequence-level auxiliary loss to prevent severe token imbalance.
However, these techniques primarily focus on training and fail to ensure balanced token assignments during inference. Instead, addressing token imbalance at inference often incurs additional resource costs. For example, DeepSeek-V3 DeepSeek-AI et al. (2024b) mitigates this issue by duplicating high-load experts and deploying them redundantly. In contrast, our approach effectively balances token assignments without introducing additional computational overhead.
3 Background and Motivation
3.1 Extremely Imbalanced Expert Utilization
A Mixture of Experts (MoE) layer consists of a collection of $n$ experts, $\{\boldsymbol{E}_{1},\boldsymbol{E}_{2},\dots,\boldsymbol{E}_{n}\}$ and a router $\boldsymbol{G}$ that dynamically selects the most relevant experts for a given input $\boldsymbol{x}$. The router computes selection scores $\boldsymbol{G}(\boldsymbol{x})$, for all experts and selects the top $k$ experts, resulting in a sparse activation:
$\mathcal{K}=\mathrm{TopK}(\mathrm{Softmax}(\boldsymbol{G}(\boldsymbol{x})),k).$ (1)
The input $\boldsymbol{x}$ is processed by the selected experts, and their outputs are combined into a weighted sum based on the router’s scores. This process is mathematically expressed as:
$\boldsymbol{y}=\sum\nolimits_{i\in\mathcal{K}}{\boldsymbol{G}(\boldsymbol{x})_{i}\cdot\boldsymbol{E}_{i}(\boldsymbol{x})},$ (2)
where $\mathcal{K}$ denotes the indices of selected experts, $\boldsymbol{G}(\boldsymbol{x})_{i}$ represents the selection score for the $i$-th expert, and $\boldsymbol{E}_{i}(\boldsymbol{x})$ is the output from the $i$-th expert. In transformer models, the MoE layer usually replaces the feed-forward network (FFN) and only activates a subset of experts for each input.
While experts in MoE can be deployed in parallel, imbalanced token-to-expert assignments lead to varying levels of expert utilization and introduce potential latency. Despite the incorporation of balancing techniques during training, the load imbalance persists during inference. To further investigate this issue, we conduct preliminary experiments to analyze expert-specific utilization patterns and assess the impact of imbalance on practical latency.
To quantify expert utilization, we measure the load across different experts. Given an input batch $\mathbf{x}\in\mathbb{R}^{b\times s\times d}$ with batch size $b$ and sequence length $s$, the total number of tokens is $t=bs$. Since each token selects $k$ out of $n$ experts, the expected token count per expert is:
$\bar{N}=\frac{tk}{n}.$ (3)
However, due to imbalanced token assignments, some experts may receive more or fewer tokens than the expected value.
Figure 2: Expert-wise load, where each load value is divided by $\bar{N}$ for clarity. To ensure generality, we visualize loads across different datasets.
Figure 2 illustrates the normalized peak tokens load for each expert to accommodate all tokens within a single layer of OLMoE, where some experts receive an excessively large number of tokens (e.g., more than seven times of average load), which resulting in significant latency. A detailed layer-by-layer analysis is provided in Appendix B.
3.2 Motivation – the Straggler Effect
Under the expert parallelism scenario, where the number of assigned tokens dictates the processing time of each expert, high-load experts become the bottleneck for overall latency within an MoE layer. Specifically, low-load experts remain idle while waiting for high-load experts to complete, leading to synchronization delays. Therefore, the latency of an MoE layer is given by:
$L\propto max(\{N_{i}\}_{i=1}^{n}),$ (4)
where $N_{i}$ represents the number of tokens assigned to the $i$-th expert, with the total token allocation satisfying $\sum\nolimits_{i=1}^{n}N_{i}=tk$. According to Eq. 4, the latency follows the Straggler Effect: the most burdened expert dictates the overall latency of the MoE layer. In the worst case, all tokens are assigned to the same group of experts, underutilizing the parallel processing capability of MoE. Conversely, distributing tokens evenly across experts maximizes computational efficiency and fully leverages the parallelism of multiple experts. With the bounds of the ideal and worst cases, the range of the highest load is given by:
$\text{max}(\{N_{i}\}_{i=1}^{n})\in[\bar{N},\frac{{}n\bar{N}}{k}].$ (5)
However, existing MoE models often adopt a dropless strategy during inference, which fails to address token imbalance and can lead to significantly increased latency.
Given that the imbalance stems from excessively high- and low-load experts, we address this issue by exploring the following questions: (1) Are there redundant tokens assigned to High-Load Experts? That is, can the MoE model maintain its performance without these excess tokens? (2) Do Low-Load Experts matter? Should these experts be removed, or should their utilization be increased to sustain or even enhance overall efficacy?
4 Methodology
(a) Capacity-Aware Token Drop
(b) Capacity-Aware Token Reroute
Figure 3: Illustrations of (a) Capacity-Aware Token Drop and (b) Capacity-Aware Token Reroute. Token Drop mitigates overload by discarding excess tokens from overloaded experts, while Token Reroute redistributes these tokens to underutilized experts, enhancing load balancing.
Token Drop Regulates the Latency of High-Load Experts
To address the question about overloaded experts, we first regulate their maximum utilization. Specifically, we introduce expert capacity to control token allocation. Given a capacity factor $\gamma$, the maximum number of tokens assigned to each expert (i.e., expert capacity) is defined as:
$C=\gamma\bar{N}.$ (6)
A higher $\gamma$ allows more tokens to be retained, but experts handling excessive tokens may introduce latency. Conversely, a lower $\gamma$ enforces stricter capacity limits, reducing latency by discarding more tokens, but at the risk of performance degradation. With the involvement of expert capacity $\gamma$, we constrain the upper bound of latency as follows:
$\max(\{N_{i}\}_{i=1}^{n})=\begin{cases}\gamma\bar{N}&\!\!\!\gamma<1\\ \text{within }[\bar{N},\gamma\bar{N}]&\!\!\!\gamma\geq 1\end{cases},$ (7)
where $\gamma$ is typically much smaller than $\frac{n}{k}$. This constraint ensures that no expert exceeds the specified capacity limit, effectively mitigating severe load imbalances and reducing latency.
Specifically, when a capacity constraint is imposed on each expert, experts must evaluate the volume of assigned tokens before execution. For experts with a load below the predefined capacity, there is no difference between capacity-constrained inference and traditional inference. However, when the load exceeds the capacity, experts must discard excess tokens to adhere to the constraint. To address this, we introduce a scoring function $\mathcal{S}$ to evaluate each token:
$\mathcal{S}(\boldsymbol{x})=\begin{bmatrix}s_{11}&s_{12}&\dots&s_{1n}\\ s_{21}&s_{22}&\dots&s_{2n}\\ \vdots&\vdots&\vdots&\vdots\\ s_{t1}&s_{t2}&\dots&s_{tn}\\ \end{bmatrix},$ (8)
where $s_{ij}$ denotes the importance score of the mapping from the $i$-th token to the $j$-th expert. With this score, each overflowed expert selectively discards those with lower scores. For the $j$-th expert:
$K_{j}=N_{j}-C,$ (9)
$\tau_{j}=\text{KthValue}(\mathcal{S}[:,j],K_{j}),$ (10)
where $K$ denotes the number of overflowed tokens, $\tau$ represents the $K$-th smallest value in $\mathcal{S}[:,j]$, serving as a threshold to filter out excess tokens:
$T_{j}\leftarrow\{t\mid\mathcal{S}[t,j]\leq\tau_{j}\}$ (11)
$\mathcal{S}[T_{j},j]\leftarrow 0,$ (12)
where $T_{j}$ denotes the indices of the rejected tokens from the $j$-th Expert. We mask the scores of the rejected tokens and prevent them from being sent to their corresponding overflowed experts.
Regarding the specific scoring function, we explore multiple efficient metrics and summarize them as follows:
Order: Discarding later tokens once earlier tokens have filled the expert capacity. This strategy was first introduced in Switch-Transformer Fedus et al. (2022) during training, and we extend it to the inference phase.
Reverse Order: Instead of discarding later tokens, this approach removes earlier tokens to comply with the expert capacity constraint.
Random: Dropping Excess tokens randomly to meet the predefined expert capacity constraints.
Score: Using the gating score $\boldsymbol{G}(\boldsymbol{x})$ as an importance indicator and discarding tokens.
Among these metrics, “Order” and “Reverse Order” are unstable, as shuffling sequences within a batch may result in different tokens being dropped Hayes et al. (2024). “Random” assumes all tokens have an equal probability of being dropped. In contrast, “Score” is stable, unaffected by sequence order within a batch.
Algorithm 1 Capacity-Aware Token Reroute
1:Expert Capacity $C$, Rounds $R$
2:Updated Scores $\mathcal{S}$
3:for $r=1,\dots,R$ do
4:     $\boldsymbol{M}\leftarrow\text{Top-$k$}(\mathcal{S},k)$
5:     $\triangleright$ Indicator for top-$k$ mappings
6:     $\mathcal{S}_{r}\leftarrow\mathcal{S}\odot\boldsymbol{M}$
7:     $\triangleright$ Mask out the non-selected
8:     for $j=1,\dots,n$ do
9:         $N_{j}\leftarrow||\mathcal{S}_{r}[:,j]||_{0}$
10:         $\triangleright$ Compute the load
11:         $K_{j}\leftarrow N_{j}-C$
12:         $\triangleright$ Number of excess tokens
13:         if $k>0$ then
14:              $\tau_{j}\leftarrow\text{KthValue}(\mathcal{S}_{r}[:,j],K_{j})$
15:              $\triangleright$ Threshold for filtering
16:              $T_{j}\leftarrow\{t\mid\mathcal{S}_{r}[t,j]\leq\tau_{j}\}$
17:              $\triangleright$ Indices of retained tokens
18:              $\mathcal{S}[T_{j},j]\leftarrow 0$
19:              $\triangleright$ Discard overflowed assignments
20:         end if
21:     end for
22:end for
Token Reroute Enhances the Utilization of Low-load Experts
Token Drop exclusively targets overloaded experts by discarding overflowed tokens that exceed expert capacity but does not address the underutilization of low-load experts. Next, we introduce Token Reroute to ensure a more balanced token-to-expert allocation. For tokens rejected by overflowed experts, we reselect underutilized experts based on the updated importance scores following Token Drop. As shown in Equation 12, masking the mappings to overflowed experts prevents these tokens from being reassigned to them, thereby encouraging their redistribution to other available experts. As illustrated in Algorithm 1, the process of masking and recomputing can be repeated iteratively, gradually achieving a more balanced token-to-expert assignment. Notably, the extra cost of Token Routing is minimal, as it occurs before tokens are assigned to each expert and operates solely on the lightweight importance matrix. Additionally, overflowed experts and tokens that have already selected $k$ experts are excluded from subsequent iterations, further minimizing the additional computational cost of the Top-$k$ operation.
5 Experiments
In this section, we conduct experiments under capacity-aware inference for MoE, with deployment details provided in Appendix A.
Table 1: Performance comparison across different capacity factors and selection metrics (i.e., Order, Reverse Order, Random, and Score). The baseline operates without capacity constraints, represented as +$\infty$+ ∞. Due to inherent randomness, we report the average performance over multiple random seeds.
Method $\gamma$ OBQA PIQA RTE WinoGrande BoolQ ARC-C HellaSwag MMLU Avg.
Baseline +$\infty$+ ∞ 45.6 80.1 53.7 71.2 74.7 54.5 79.4 52.5 64.0
Order 3.0 43.6 75.8 53.1 71.2 74.4 50.5 78.6 51.7 62.4
Reverse Order 44.8 76.1 50.9 71.0 74.8 52.8 78.7 51.7 62.6
Random 44.0 76.3 52.0 71.3 75.0 53.5 78.9 51.8 62.9
Score 45.2 79.9 51.6 71.7 74.4 55.2 79.2 51.9 63.6
Order 2.0 42.0 71.5 53.1 71.2 74.2 49.5 76.6 48.4 60.8
Reverse Order 41.8 71.8 52.7 71.0 73.9 49.4 76.4 49.2 60.8
Random 41.2 75.2 52.7 71.0 74.1 50.1 76.8 49.4 61.3
Score 44.0 79.7 53.8 72.1 74.2 54.4 78.4 50.4 63.4
Order 1.5 38.8 67.1 48.7 68.5 73.3 46.3 54.0 43.7 55.1
Reverse Order 40.2 67.3 52.7 70.1 72.7 45.5 54.4 45.2 56.0
Random 39.6 72.1 57.8 68.3 73.8 45.8 74.2 45.2 59.6
Score 43.2 76.1 56.3 69.9 73.4 52.9 77.1 47.5 62.1
Order 1.0 36.0 60.2 53.4 62.6 69.6 38.7 58.0 36.9 51.9
Reverse Order 36.2 59.5 50.5 63.3 69.4 39.4 58.7 38.7 52.0
Random 34.0 63.1 55.2 60.8 70.2 40.5 66.9 35.7 53.3
Score 40.4 71.5 57.0 64.3 71.9 45.4 72.0 39.7 57.8
5.1 Token Drop for High-load Experts
Investigation on Token Drop Metrics
To assess the effectiveness of different metrics in regulating token load to the target capacity, we compare various approaches on OLMoE by discarding excess tokens and applying a range of capacity factors. As shown in Table 1, varying the dropping metrics impacts performance at different levels. With higher capacities, the model maintains comparable performance even when using naive selection methods like “Random”. However, as the capacity factor decreases, performance degradation becomes more pronounced, particularly for “Order”, “Reverse Order”, and “Random”. Notably, “Score” consistently outperforms other methods by a large margin, demonstrating the effectiveness of leveraging gating scores as an importance measure. Consequently, we adopt "Score" as the default metric.
(a) Single MoE Layer
(b) End-to-End Model
Figure 4: Speedup achieved by capacity-aware inference compared to the baseline without capacity constraints.
Efficiency Gains from Capacity-Constrained Inference
We next explore the efficiency improvements achieved by imposing expert capacity. Specifically, we employ distributed inference using eight A30 GPUs, utilizing an 8-way Data Parallelism (DP) and 8-way Expert Parallelism (EP) strategy through the Megatron-LM framework Shoeybi et al. (2019). In Mixtral-8$\times$7B-Instruct model, each GPU hosts a single expert, whereas, in models like OLMoE-Instruct, multiple experts must be deployed on a single GPU (e.g., eight experts per GPU) due to GPU resource constraints.
As illustrated in Figure 4, imposing constraints on expert capacity considerably accelerates inference across the four tested MoE models, in comparison to the baseline model without capacity limitations. The enhanced efficiency of each MoE layer (Figure 4 (a)) contributes to faster end-to-end inference (Figure 4 (b)). For instance, when the capacity factor $\gamma$ is set to 1.5, this configuration achieves a 1.94$\times$ speedup for a single MoE module and a 1.37$\times$ speedup for the entire end-to-end model. Furthermore, it is evident that the efficiency improves as the capacity decreases.
Notably, the degree of acceleration is influenced by the numerical relationship between the number of experts and GPUs in Expert Parallelism. In Mixtral-8×8$\times$8 ×7B-Instruct, one-to-one deployment maximizes the effectiveness of capacity-aware inference. Conversely, in models where multiple experts share a single GPU, the acceleration gains are relatively modest. This occurs because the aggregated load from multiple experts diminishes the proportion of reduced load, which is achieved by limiting the straggler expert. Additionally, we believe that further system optimizations or allocating more GPUs for expert distribution would further enhance the potential of capacity-aware inference.
Figure 5: Analysis of dropped tokens with respect to capacity factors.
Mitigating the Straggler Effect with Minimal Token Discarding
Given that expert capacity enforces MoE layers to discard overflowed tokens, we next establish the relationship between expert capacity and the corresponding number of dropped tokens. Specifically, for a capacity factor $\gamma$, the total proportion of dropped tokens is given by:
$DT=\frac{\sum_{i=1}^{n}\text{ReLU}(N_{i}-\gamma\bar{N})}{\sum_{i=1}^{n}N_{i}},$ (13)
where $\text{ReLU}(N_{i}-\gamma\bar{N})$ represent the number of dropped tokens for the $i$-th expert.
Figure 5 visualizes the number of dropped tokens across different capacity factors for various test datasets, with a more detailed illustration provided in Appendix C. Although the most overloaded expert receives much more tokens than the expected number of tokens $\bar{N}$, regulating the maximum capacity has a limited impact on the overall number of accommodated tokens (e.g., reducing it by less than 20% under a capacity factor of 2.0), thereby maintaining competitive performance even after discarding overflow tokens. Moreover, dropping a small proportion of overflowed tokens can significantly reduce the latency caused by overloaded experts (e.g., dropping 12% of overloaded tokens promotes the inference speed by 94% in Mixtral-8$\times$7B-Instruct), highlighting the effectiveness of capacity-constrained inference in improving both performance and efficiency.
Table 2: Comparison of Expert Drop, Token Drop and Token Reroute. The capacity factor $\gamma$ is set to 2.0 for OLMoE and DeepSeek-V2-Lite, and 1.5 for Qwen1.5-MoE-Chat and Mixtral-8×7B-Instruct. In Expert Drop, we only skip one out of eight experts for Mixtral-8$\times$7B-Instruct, and 10% of the lowest load experts for other models.
Model Method OBQA PIQA RTE WinoGrande BoolQ ARC-C HellaSwag MMLU GSM8K Avg.
OLMoE-Instruct Baseline 47.6 80.2 67.9 69.9 80.7 57.0 80.6 52.8 35.1 63.5
Expert Drop 44.6 76.9 64.0 67.6 78.2 54.4 77.0 50.6 31.6 60.5
Token Drop 47.8 77.9 64.6 69.2 80.0 57.2 79.7 51.5 32.4 62.3
Token Reroute 47.2 79.4 66.3 70.5 80.9 57.1 80.3 52.3 34.4 63.2
Qwen1.5-MoE-Chat Baseline 42.4 79.9 72.9 70.0 81.3 54.1 80.4 59.8 52.0 65.9
Expert Drop 41.4 78.7 71.2 68.6 80.6 52.9 79.1 58.1 49.4 64.4
Token Drop 40.4 78.8 72.6 69.1 80.9 53.0 80.0 59.3 51.9 65.1
Token Reroute 43.4 79.1 72.6 69.6 81.1 53.4 80.3 59.3 52.1 65.6
DeepSeek-V2-Lite-Chat Baseline 45.4 81.4 72.6 75.5 82.9 61.0 81.5 57.3 66.4 69.3
Expert Drop 41.8 77.6 71.9 72.5 81.6 57.1 75.5 53.3 56.0 65.3
Token Drop 45.2 78.3 72.6 74.0 83.2 59.3 80.9 57.3 62.7 68.2
Token Reroute 45.4 79.4 73.3 75.4 83.2 60.4 81.5 57.2 64.1 68.9
Mixtral-8$\times$7B-Instruct Baseline 47.4 84.8 71.8 82.5 88.5 71.7 87.5 70.2 64.2 74.3
Expert Drop 46.8 83.2 70.1 81.3 87.6 67.1 85.6 66.2 62.3 72.2
Token Drop 46.4 83.3 71.7 82.2 88.3 71.2 87.4 69.1 64.7 73.8
Token Reroute 47.8 85.0 71.8 83.0 88.6 71.5 87.6 70.2 64.6 74.5
5.2 Token Reroute to Low-load Experts
Besides the experts overloaded with tokens, some low-load experts receive only a few tokens, raising important questions: Are these low-load experts redundant and removable, or should they be leveraged to balance token allocation? Recent works Lu et al. (2024); He et al. (2024) remove less important experts to improve efficiency, while our proposed Token Reroute increases their utilization by redistributing tokens for a more balanced assignment. Next, we investigate the significance of low-load experts and validate the effectiveness of Token Reroute.
The Critical Role of Low-Load Experts
To explore the impact of low-load experts, we further compare dropping tokens (i.e., Token Drop) with skipping experts (i.e., Expert Drop). For Expert Drop, we adopt a conservative strategy that dynamically skips the 10% of experts with the lowest token loads. Notably, the proportion of tokens removed in Expert Drop is significantly lower than in Token Drop (2% in Expert Drop vs. 12% in Token Drop on OLMoE-Instruct).
Despite this, as shown in Table 2, Expert Drop experiences significant performance degradation and is outperformed by Token Drop by a large margin. Moreover, due to the small proportion of tokens assigned to low-load experts, removing these experts provides only marginal improvements in inference speed (less than a 5% speedup). These findings indicate that retaining low-load experts better preserves the performance of MoE models.
Effectiveness of Token Reroute
We examine the effectiveness of utilizing low-load experts by rerouting overflowed tokens to them (i.e., Token Reroute) instead of simply discarding these tokens to meet the target capacity. Comparing Token Reroute with Token Drop, redistributing excess tokens to low-load experts enhances performance, yielding a 0.9% improvement in the average performance of Qwen1.5-MoE-Chat. Furthermore, considering the performance degradation observed in Expert Drop, our findings highlight the crucial role of low-load experts in maintaining model effectiveness.
(a) OLMoE
(b) OLMoE-Instruct
(c) Qwen1.5-MoE
(d) Qwen1.5-MoE-Chat
Figure 6: Performance change as capacity factors decrease from 3.0 to 0.0.
5.3 Ablation Study
Model-Specific Imbalanced Property
We explore the imbalance property in various models, such as OLMoE, DeepSeek-V2-Lite and Qwen1.5-MoE, which differ in both architecture (e.g., depth and width) and training strategies (e.g., training from scratch Muennighoff et al. (2024); DeepSeek-AI et al. (2024a) vs. training after upcycling Jiang et al. (2024); Team (2024b)).
On the one hand, our findings in Appendix B reveal different training strategies result in significantly varying levels of imbalance. Specifically, MoE models trained from scratch exhibit a much higher degree of imbalance. For instance, OLMoE and DeepSeek-V2-Lite experience peak expert-wise token allocations exceeding $5\bar{N}$, whereas Qwen1.5-MoE and Mixtral are upcycled from dense language models, maintain a more balanced distribution, with peak expert-wise allocations staying below $3\bar{N}$. This is because upcycling initializes experts with the same parameters Komatsuzaki et al. (2023), while experts are randomly initialized when training from scratch, leading to a greater imbalance in token assignments Lo et al. (2024).
On the other hand, despite the widespread use of auxiliary balance loss in MoE training, it does not guarantee balanced token assignments across experts, as token distribution still varies significantly during inference on test data. This necessitates integrating expert capacity into the inference process.
Capacity Factor
Beyond the specific capacity values presented in Table 1, we further investigate a wide range of capacity factors in Figure 6, spanning from 0.0 to 3.0. We exclude values exceeding 3.0, as their performance closely aligns with capacity-agnostic scenarios. By analyzing the performance changes when decreasing the capacity factor, we find that setting $\gamma$ to 2.0 is sufficient to maintain performance comparable to the original models. This means allowing each expert to handle utmost twice the average number of tokens maintains the performance of capacity-free inference. However, maintaining the performance becomes challenging under low capacity factors (e.g., 1.0), as high-load experts experience significant token drops.
Figure 7: Ablation study on the number of iterations for Token Reroute.
Rerouting Round
The Token Reroute Algorithm iteratively discourages tokens from selecting overflowed experts while encouraging their rerouting to other available experts, promoting a more balanced token distribution. As iterations progress, lower-score token-to-expert mappings are increasingly considered, further refining the redistribution process. Figure 7 illustrates the results across different iterations. In the first iteration, only excess tokens are dropped, while token rerouting becomes more prominent in subsequent rounds. The first two rounds of token rerouting lead to significant performance improvements, after which performance begins to saturate. We attribute this saturation to the selection of additional low-score token-to-expert mappings, which have a smaller impact compared to the dominant token assignments.
6 Conclusion
In this paper, we first identify the imbalanced token-to-expert assignment in Mixture of Experts (MoE) and introduce the Straggler Effect in MoE inference, where the highest-load expert becomes the efficiency bottleneck, dictating overall latency. To address this issue, we propose Capacity-Aware Token Drop, which alleviates excessive loads on individual experts, and Capacity-Aware Token Reroute, which enhances the utilization of underutilized experts. Our findings and proposed methods offer valuable insights and effective strategies for improving MoE inference efficiency.
Limitations
Despite the progress we have made, our work still has certain limitations. First, our study focuses on a subset of MoE models, including OLMoE, Qwen1.5-MoE, DeepSeek-V2-Lite, and Mixtral. A promising direction for future work is to extend our analysis and methods to other architectures to further validate their effectiveness across a broader range of MoE models. Second, our proposed Capacity-Aware Token Drop and Capacity-Aware Token Reroute strategies could be integrated into the training process, potentially enhancing both performance and efficiency.
Appendix A Implementation Details
Models
We mainly focus on lightweight MoE models (less than 20B parameter budget). We conduct experiments on OLMoE Muennighoff et al. (2024), Qwen1.5-MoE Team (2024b), DeepSeek-V2-Lite DeepSeek-AI et al. (2024a) and Mixtral Jiang et al. (2024), due to their competitive performance and widespread adoption.
Datasets
To evaluate model performance, we report normalized zero-shot or few-shot accuracy on the LM-Harness benchmark. The number of shots for each task is detailed in Table 3, which includes multiple tasks: ARC-C Clark et al. (2018), BoolQ Clark et al. (2019), HellaSwag Zellers et al. (2019), MMLU Hendrycks et al. (2021), OBQA Mihaylov et al. (2018), PIQA Bisk et al. (2019), RTE Wang et al. (2019), WinoGrande ai2 (2019) and GSM8K Cobbe et al. (2021). The evaluation code is based on EleutherAI’s LM Harness framework Gao et al. (2023).
Table 3: Experimental settings for evaluation tasks. “Norm” refers to the normalization performed with respect to the length of the input.
 Task  Number of few-shot  Metric
BoolQ 0 Accuracy
RTE 0 Accuracy
OBQA 0 Accuracy (Norm)
PIQA 0 Accuracy (Norm)
MMLU 5 Accuracy
WinoGrande 5 Accuracy
GSM8K 5 Exact Match
HellaSwag 10 Accuracy (Norm)
ARC-C 25 Accuracy (Norm)
Appendix B Layer-wise Expert Load
To analyze imbalanced token assignments, we measure the expert load for each expert by tracking the peak expert load while running MoE models on various test datasets. Figure 8, 9, 10 and 11 present the full results for the normalized layer-wise expert load for OLMoE, DeepSeek-V2, Qwen1.5-MoE, and Mixtral-8$\times$7B-Instruct, respectively.
Figure 8: Layer-wise expert load in OLMoE-Instruct.
Figure 9: Layer-wise expert load in Deepseek-V2-Lite.
Figure 10: Layer-wise expert load in Qwen1.5-MoE-Chat.
Figure 11: Layer-wise expert load in Mixtral-8$\times$7B-Instruct.
Figure 12: Dropped tokens with respect to capacity factors in OLMoE-Instruct.
Figure 13: Dropped tokens with respect to capacity factors in DeepSeek-V2-Chat.
Figure 14: Dropped tokens with respect to capacity factors in Qwen-1.5-MoE-Chat.
Figure 15: Dropped tokens with respect to capacity factors in Mixtral-8$\times$7B-Instruct.
Appendix C Dropped tokens Calculation
Based on Equation 13, we calculate the total number of dropped tokens across experts in each layer under different capacity factors, as illustrated in Figures 12, 14, 13, and 15.

### v3

@@
 Abstract
-The Mixture of Experts (MoE) is an effective architecture for scaling large language models by leveraging sparse expert activation, optimizing the trade-off between performance and efficiency.
-However, under expert parallelism, MoE suffers from inference inefficiencies due to imbalanced token-to-expert assignment, where some experts are overloaded while others remain underutilized.
-This imbalance leads to poor resource utilization and increased latency, as the most burdened expert dictates the overall delay, a phenomenon we define as the Straggler Effect.
-To mitigate this, we propose Capacity-Aware Inference, including two key techniques: (1) Capacity-Aware Token Drop, which discards overloaded tokens to regulate the maximum latency of MoE, and (2) Capacity-Aware Token Reroute, which reallocates overflowed tokens to underutilized experts, balancing the token distribution.
-These techniques collectively optimize both high-load and low-load expert utilization, leading to a more efficient MoE inference pipeline.
-Extensive experiments demonstrate the effectiveness of our methods, showing significant improvements in inference efficiency, e.g., 0.2% average performance increase and a 1.94$\times$ inference speedup on Mixtral-8$\times$7B-Instruct.
-Capacity-Aware Inference: Mitigating the Straggler Effect in
-Mixture of Experts
-Shwai He1 Weilin Cai2 Jiayi Huang2 Ang Li1 1University of Maryland, College Park 2The Hong Kong University of Science and Technology (Guangzhou) shwaihe@umd.edu, angliece@umd.edu
+The Mixture of Experts (MoE) is an effective architecture for scaling large language models by leveraging sparse expert activation to balance performance and efficiency.
+However, under expert parallelism, MoE suffers from inference inefficiencies due to imbalanced token-to-expert assignment, where underloaded experts complete computations early but must wait for overloaded experts, leading to global delays.
+We define this phenomenon as the Straggler Effect, as the most burdened experts dictate the overall inference latency.
+To address this, we first propose Capacity-Aware Token Drop, which enforces expert capacity limits by discarding excess tokens from overloaded experts, effectively reducing load imbalance with minimal performance impact (e.g., $30\%$ speedup with only $0.9\%$ degradation on OLMoE).
+Next, given the presence of low-load experts remaining well below the capacity threshold, we introduce Capacity-Aware Expanded Drop, which allows tokens to include additional local experts in their candidate set before enforcing strict local capacity constraints, thereby improving load balance and enhancing the utilization of underused experts.
+Extensive experiments on both language and multimodal MoE models demonstrate the effectiveness of our approach, yielding substantial gains in expert utilization, model performance, and inference efficiency, e.g., applying Expanded Drop to Mixtral-8$\times$7B-Instruct yields a 0.2% average performance improvement and a 1.85$\times$ inference speedup.
 1 Introduction
-In recent years, the rapid evolution of Large Language Models (LLMs) OpenAI (2024); Team (2024a); DeepSeek-AI et al.
-(2024b) has driven a wave of innovations, continuously expanding the frontiers of AI research and applications.
+Figure 1: Illustration of the Straggler Effect in MoE Inference, where the most burdened experts dictate the overall latency.
+In recent years, the rapid evolution of Large Language Models (LLMs) [32, 38, 11] has driven a wave of innovations, continuously expanding the frontiers of AI research and applications.
 Among the model architectural innovations, the Mixture of Experts (MoE) framework has emerged as a pivotal technique for optimizing the cost-performance trade-off in LLMs.
-Specifically, MoE Shazeer et al.
-(2017a); Fedus et al.
-(2022) enhances scalability by integrating multiple experts while activating only a subset per input.
+Specifically, MoE [35, 14] enhances scalability by integrating multiple experts while activating only a subset per input.
 This selective activation substantially improves model performance without a corresponding increase in computational cost, effectively balancing efficiency and performance.
-Despite the success of MoE, a key efficiency challenge lies in the imbalanced token-to-expert distribution, which results in some experts being overloaded while others remain underutilized Lepikhin et al.
-(2021); Zoph et al.
-(2022).
-In the context of expert parallelism, low-load experts complete their computations faster but must wait for high-load experts to finish, leading to inefficient resource utilization and increased latency.
-As illustrated in Figure 1, this phenomenon is defined as the Straggler Effect, where the most burdened expert dictates the overall latency of the MoE layer.
-Figure 1: Illustration of the Straggler Effect in MoE Inference, where the most burdened expert (i.e., Straggler Expert) dictates the overall latency.
-While auxiliary balance losses have been incorporated into the training process to alleviate imbalance Shazeer et al.
-(2017a); Fedus et al.
-(2022); DeepSeek-AI et al.
-(2024b), these techniques remain ineffective in mitigating imbalance during inference.
+Despite the success of MoE, a key efficiency challenge lies in the imbalanced token-to-expert distribution, which results in some experts being overloaded while others remain underutilized [26, 46].
+In distributed GPU settings, experts are typically sharded across multiple devices, with each GPU responsible for a subset of the experts.
+Under expert parallelism, low-load experts complete their computations earlier but must wait for overloaded experts to finish, as synchronization barriers are required before proceeding to the next stage.
+This expert-level straggler effect further propagates to device-level delays, where GPUs hosting lighter expert workloads are stalled by GPUs hosting heavier workloads, leading to inefficient resource utilization and increased end-to-end latency during inference.
+As illustrated in Figure 1, this phenomenon is referred to as the Straggler Effect, where the heavily loaded experts determine the overall latency of imbalanced MoE inference.
+While auxiliary balance losses have been incorporated into the training process to alleviate imbalance [35, 14, 11], these techniques remain ineffective in mitigating imbalance during inference.
 Specifically, as shown in Figure 1, our findings reveal a highly uneven token distribution among experts, with the highest-load expert handling more than seven times the expected average load.
-Moreover, managing such an imbalance during inference often incurs additional resource costs.
-For instance, DeepSeek-V3 addresses this issue by duplicating high-load experts and deploying them redundantly DeepSeek-AI et al.
-(2024b).
-This motivates us to explore efficient token-to-expert assignment through two key questions: (1) How to prevent extreme overloading in high-load experts?
-and (2) How can we utilize the available capacity of low-load experts?
-We address these challenges through Capacity-Aware Inference.
-For high-load experts, we introduce Capacity-Aware Token Drop, a technique that removes excessive tokens from overloaded experts.
-On the one hand, eliminating these excess tokens alleviates extreme load imbalances, significantly enhancing efficiency.
-On the other hand, dropping these tokens minimally impacts model performance, as they constitute only a small fraction of the total tokens.
-For low-load experts, we extend Token Drop with Capacity-Aware Token Reroute, which leverages the available capacity of underutilized experts to process overflowed tokens.
-Token Reroute further enhances the performance of Token Drop within the capacity-constrained inference framework.
-Extensive experimental results validate the effectiveness of our proposed techniques, demonstrating significant improvements in both efficiency and performance, e.g., 0.2% average performance increase and a 1.94$\times$ inference speedup in Mixtral-8$\times$7B-Instruct.
-In short, our contributions are in three folds:
+Moreover, managing such imbalance during inference often incurs additional resource overhead.
+For example, DeepSeek-V3 mitigates this issue by duplicating high-load experts and deploying them redundantly across devices [11].
+This motivates us to explore efficient token-to-expert assignment by addressing the key question: How can we prevent extreme overloading of heavily utilized experts?
+We propose Capacity-Aware Inference to address this challenge.
+Specifically, for high-load experts, we introduce Capacity-Aware Token Drop, which imposes a maximum capacity constraint and discards excess tokens from overloaded experts.
+This approach alleviates severe load imbalance and significantly improves efficiency, while maintaining model performance since the dropped tokens represent only a small fraction of the total workload, e.g., OLMoE achieves a 30% speedup in MoE layers with just a 0.9% performance degradation.
+After removing excess tokens from overloaded experts, we observe that some low-load experts remain significantly underutilized relative to the predefined capacity constraints, yet must still wait for other experts to complete their computations.
+This leads us to a second key question: How can we effectively leverage the available capacity of underutilized experts?
+For low-load experts, we extend Token Drop with Capacity-Aware Expanded Drop, which further utilizes the available capacity of underutilized experts to handle overflow tokens from high-load experts.
+Specifically, under expert parallelism across multiple GPUs, Expanded Drop allows each token to consider additional candidate experts on the same device while still enforcing strict local capacity constraints.
+This expanded selection improves the utilization of low-load experts and enhances the representational capacity of MoE models under capacity-constrained scenarios.
+Extensive experimental results validate the effectiveness of our proposed techniques, demonstrating significant improvements in both efficiency and performance, e.g., applying Expanded Drop to Mixtral-8$\times$7B-Instruct yields a 0.2% average performance improvement and a 1.85$\times$ inference speedup.
+Moreover, in multimodal models, we identify redundancy among image tokens and show that applying aggressive capacity constraints (e.g., setting the maximum to half of the average expert load) can still maintain performance.
+In short, our contributions are in four folds:
   • We identify the Straggler Effect caused by token imbalance at inference time in Mixture of Experts, highlighting the optimization potential for reducing latency.
-  • Toward token imbalance, we propose Token Drop and Token Reroute to enhance the balanced utilization of experts.
-  • Experimental results demonstrate the effectiveness of Token Drop and Token Reroute, achieving significant improvements in expert utilization and inference efficiency.
+  • Toward the high-load experts, we propose Capacity-Aware Token Drop, which enforces capacity constraints by discarding excess tokens assigned to overloaded experts, thereby mitigating extreme load imbalance.
+  • To better utilize underloaded experts, we introduce Capacity-Aware Expanded Drop, which expands the candidate expert set to include additional local experts, further improving load balance and model performance.
+  • Extensive experiments on both language and multimodal models validate the effectiveness of our approach, demonstrating substantial improvements in inference efficiency with comparable performance.
 2 Related Works
 Mixture of Experts Models
-The Mixture of Experts (MoE) is a kind of neural network architecture with an extended set of parameters (referred to as “experts”) controlled by a router, which is first introduced in the context of conditional computation (Jacobs et al., 1991; Jordan and Jacobs, 1994).
-The potential of sparse activation in MoE is subsequently exploited by Shazeer et al.
-(2017b) for efficient training and inference on pretrained models with special designs, opening the door for MoE in various vision (Riquelme et al., 2021) and language (Lepikhin et al., 2020; Du et al., 2022; Fedus et al., 2022) scenarios.
-Attributed to its exceptional efficiency, MoE has been adopted as a foundational framework in the designs of large language models (Jiang et al., 2024; Dai et al., 2024; Xue et al., 2024a; Zhu et al., 2024; Team, 2024b), achieving superior scaling laws at low computational costs.
-Despite these advancements, MoE still faces efficiency challenges in both training and inference Cai et al.
-(2024), and our work specifically focuses on enhancing inference-time efficiency.
+The Mixture of Experts (MoE) is a kind of neural network architecture with an extended set of parameters (referred to as “experts”) controlled by a router, which is first introduced in the context of conditional computation [21, 23].
+The potential of sparse activation in MoE is subsequently exploited by [34] for efficient training and inference on pretrained models with special designs, opening the door for MoE in various vision [33] and language [25, 13, 14] scenarios.
+Attributed to its exceptional efficiency, MoE has been adopted as a foundational framework in the designs of large language models [22, 9, 41, 45, 39], achieving superior scaling laws at low computational costs.
+Despite these advancements, MoE still faces efficiency challenges in both training and inference [3], and our work specifically focuses on enhancing inference-time efficiency.
 Imbalance in Mixture of Experts
-The imbalance in token-to-expert assignments Zhou et al.
-(2022); Chen et al.
-(2022) poses a significant challenge to the deployment of Mixture of Experts (MoE).
-This imbalance leads to inefficiencies in memory, computation, and communication (He et al., 2023; Song et al., 2023; Xue et al., 2024b), making it a critical bottleneck for MoE scalability and deployment.
-To mitigate this issue, an auxiliary balance loss Shazeer et al.
-(2017a) is incorporated into the training process to encourage more uniform token distribution across experts.
-Additionally, various training strategies have been introduced to further balance token assignments: Switch-Transformer Fedus et al.
-(2022) and DeepSeek-V2 DeepSeek-AI et al.
-(2024a) implement Token Drop to alleviate expert overload, while DeepSeek-V3 DeepSeek-AI et al.
-(2024b) introduces an additional sequence-level auxiliary loss to prevent severe token imbalance.
+The imbalance in token-to-expert assignments [44, 5] poses a significant challenge to the deployment of MoE.
+This imbalance leads to inefficiencies in computation, communication, and memory [18, 37, 42], making it a critical bottleneck for MoE scalability and deployment.
+To mitigate this issue, an auxiliary balance loss [35] is incorporated into the training process to encourage more uniform token distribution across experts.
+Additionally, various training strategies have been introduced to further balance token assignments: Switch-Transformer [14] and DeepSeek-V2 [10] implement Token Drop to alleviate expert overload, while DeepSeek-V3 [11] introduces an additional sequence-level auxiliary loss to prevent severe token imbalance.
 However, these techniques primarily focus on training and fail to ensure balanced token assignments during inference.
 Instead, addressing token imbalance at inference often incurs additional resource costs.
-For example, DeepSeek-V3 DeepSeek-AI et al.
-(2024b) mitigates this issue by duplicating high-load experts and deploying them redundantly.
+For example, DeepSeek-V3 [11] mitigates this issue by duplicating high-load experts and deploying them redundantly.
 In contrast, our approach effectively balances token assignments without introducing additional computational overhead.
@@
 In transformer models, the MoE layer usually replaces the feed-forward network (FFN) and only activates a subset of experts for each input.
-While experts in MoE can be deployed in parallel, imbalanced token-to-expert assignments lead to varying levels of expert utilization and introduce potential latency.
+While experts in MoE models are often deployed in parallel across distributed GPUs, imbalanced token-to-expert assignments lead to varying levels of expert utilization and potential latency.
 Despite the incorporation of balancing techniques during training, the load imbalance persists during inference.
 To further investigate this issue, we conduct preliminary experiments to analyze expert-specific utilization patterns and assess the impact of imbalance on practical latency.
+Figure 2: Expert-wise load, where each load value is divided by $\bar{N}$ for clarity. To ensure generality, we visualize loads across different datasets.
 To quantify expert utilization, we measure the load across different experts.
@@
 However, due to imbalanced token assignments, some experts may receive more or fewer tokens than the expected value.
-Figure 2: Expert-wise load, where each load value is divided by $\bar{N}$ for clarity. To ensure generality, we visualize loads across different datasets.
-Figure 2 illustrates the normalized peak tokens load for each expert to accommodate all tokens within a single layer of OLMoE, where some experts receive an excessively large number of tokens (e.g., more than seven times of average load), which resulting in significant latency.
-A detailed layer-by-layer analysis is provided in Appendix B.
+Figure 2 illustrates the normalized peak tokens load for each expert to accommodate all tokens within a single layer of OLMoE, where some experts receive an excessively large number of tokens (e.g., more than seven times the average number of tokens), leading to severe load imbalance and, consequently, significant latency.
+A detailed layer-by-layer analysis is provided in Appendix E.
 3.2 Motivation – the Straggler Effect
@@
 According to Eq.
-4, the latency follows the Straggler Effect: the most burdened expert dictates the overall latency of the MoE layer.
+4, the latency follows the Straggler Effect: the most burdened experts dictate the overall latency of the MoE layer.
 In the worst case, all tokens are assigned to the same group of experts, underutilizing the parallel processing capability of MoE.
@@
 However, existing MoE models often adopt a dropless strategy during inference, which fails to address token imbalance and can lead to significantly increased latency.
-Given that the imbalance stems from excessively high- and low-load experts, we address this issue by exploring the following questions: (1) Are there redundant tokens assigned to High-Load Experts?
-That is, can the MoE model maintain its performance without these excess tokens?
-(2) Do Low-Load Experts matter?
-Should these experts be removed, or should their utilization be increased to sustain or even enhance overall efficacy?
+Given that the imbalance stems from excessively high- and low-load experts, we address this issue by exploring the following questions: (1) For high-load experts, are there redundant tokens that can be dropped without causing significant performance degradation?
+(2) For low-load experts that must wait for high-load experts to complete forward passes, is there an opportunity to enhance their utilization and improve performance without incurring substantial additional cost?
 4 Methodology
-(a) Capacity-Aware Token Drop
-(b) Capacity-Aware Token Reroute
-Figure 3: Illustrations of (a) Capacity-Aware Token Drop and (b) Capacity-Aware Token Reroute.
-Token Drop mitigates overload by discarding excess tokens from overloaded experts, while Token Reroute redistributes these tokens to underutilized experts, enhancing load balancing.
 Token Drop Regulates the Latency of High-Load Experts
@@
 This constraint ensures that no expert exceeds the specified capacity limit, effectively mitigating severe load imbalances and reducing latency.
+Note that tokens are distributed across devices under expert and data parallelism.
+To avoid additional communication overhead, we apply capacity constraints to tokens within each local device, similar to the constraints used during training [14].
+This ensures that all devices respect the limits, maintaining strict control over token flow to the experts.
 Specifically, when a capacity constraint is imposed on each expert, experts must evaluate the volume of assigned tokens before execution.
@@
 With this score, each overflowed expert selectively discards those with lower scores.
-For the $j$-th expert:
-$K_{j}=N_{j}-C,$ (9)
-$\tau_{j}=\text{KthValue}(\mathcal{S}[:,j],K_{j}),$ (10)
-where $K$ denotes the number of overflowed tokens, $\tau$ represents the $K$-th smallest value in $\mathcal{S}[:,j]$, serving as a threshold to filter out excess tokens:
-$T_{j}\leftarrow\{t\mid\mathcal{S}[t,j]\leq\tau_{j}\}$ (11)
-$\mathcal{S}[T_{j},j]\leftarrow 0,$ (12)
-where $T_{j}$ denotes the indices of the rejected tokens from the $j$-th Expert. We mask the scores of the rejected tokens and prevent them from being sent to their corresponding overflowed experts.
-Regarding the specific scoring function, we explore multiple efficient metrics and summarize them as follows:
+Let $\mathcal{J}$ be the set of overflowed experts:
+$\tau_{\mathcal{J}}=\text{KthValue}(\mathcal{S}_{\mathcal{J}},C),$ (9)
+where $\tau_{\mathcal{J}}$ represents the thresholds, i.e., $C$-th highest value in $\mathcal{S}_{\mathcal{J}}$, serving as a threshold to filter out excess tokens:
+$T_{\mathcal{J}}\leftarrow\left\{(t,j)\;\middle|\;t\in[1,\dots,N],\;j\in\mathcal{J},\;\mathcal{S}[t,j]\geq\tau_{\mathcal{J}}[j]\right\}$ (10)
+$\mathcal{S}_{\mathcal{J}}\leftarrow\mathcal{S}_{\mathcal{J}}\odot M_{\mathcal{J}},\text{where}~{}M_{\mathcal{J}}\leftarrow\mathbb{1}\left[\mathcal{S}_{\mathcal{J}}\geq\tau_{\mathcal{J}}\right],$ (11)
+where $T_{\mathcal{J}}$ denotes the token indices retained by the experts indexed in $\mathcal{J}$.
+The scores of rejected tokens are masked to prevent them from being routed to their corresponding overflowed experts.
+Regarding the specific scoring function, we explore multiple metrics and summarize them as follows:
 Order: Discarding later tokens once earlier tokens have filled the expert capacity.
-This strategy was first introduced in Switch-Transformer Fedus et al.
-(2022) during training, and we extend it to the inference phase.
+This strategy was first introduced in Switch-Transformer [14] during training, and we extend it to the inference phase.
 Reverse Order: Instead of discarding later tokens, this approach removes earlier tokens to comply with the expert capacity constraint.
@@
 Score: Using the gating score $\boldsymbol{G}(\boldsymbol{x})$ as an importance indicator and discarding tokens.
-Among these metrics, “Order” and “Reverse Order” are unstable, as shuffling sequences within a batch may result in different tokens being dropped Hayes et al.
-(2024).
+Among these metrics, “Order” and “Reverse Order” are unstable, as shuffling sequences within a batch may result in different tokens being dropped [17].
 “Random” assumes all tokens have an equal probability of being dropped.
 In contrast, “Score” is stable, unaffected by sequence order within a batch.
-Algorithm 1 Capacity-Aware Token Reroute
-1:Expert Capacity $C$, Rounds $R$
-2:Updated Scores $\mathcal{S}$
-3:for $r=1,\dots,R$ do
-4:     $\boldsymbol{M}\leftarrow\text{Top-$k$}(\mathcal{S},k)$
-5:     $\triangleright$ Indicator for top-$k$ mappings
-6:     $\mathcal{S}_{r}\leftarrow\mathcal{S}\odot\boldsymbol{M}$
-7:     $\triangleright$ Mask out the non-selected
-8:     for $j=1,\dots,n$ do
-9:         $N_{j}\leftarrow||\mathcal{S}_{r}[:,j]||_{0}$
-10:         $\triangleright$ Compute the load
-11:         $K_{j}\leftarrow N_{j}-C$
-12:         $\triangleright$ Number of excess tokens
-13:         if $k>0$ then
-14:              $\tau_{j}\leftarrow\text{KthValue}(\mathcal{S}_{r}[:,j],K_{j})$
-15:              $\triangleright$ Threshold for filtering
-16:              $T_{j}\leftarrow\{t\mid\mathcal{S}_{r}[t,j]\leq\tau_{j}\}$
-17:              $\triangleright$ Indices of retained tokens
-18:              $\mathcal{S}[T_{j},j]\leftarrow 0$
-19:              $\triangleright$ Discard overflowed assignments
-20:         end if
-21:     end for
-22:end for
-Token Reroute Enhances the Utilization of Low-load Experts
+Notably, there is virtually no additional computational overhead associated with calculating these metrics, and the dropping operation incurs minimal cost compared to the intensive computations performed by the experts.
+(a) Token Drop
+(b) Expanded Drop
+Figure 3: Illustration of Capacity-Aware Token Drop (a) and Expanded Drop (b).
+Both methods first select experts based on gating scores.
+In Token Drop, tokens exceeding the local device capacity are discarded prior to All-to-All communication.
+Expanded Drop enhances expert utilization by allowing each token to consider additional $m$ candidate experts on the same device while still enforcing strict local capacity constraints.
+Expanded Drop Enhances the Utilization of Low-load Experts
 Token Drop exclusively targets overloaded experts by discarding overflowed tokens that exceed expert capacity but does not address the underutilization of low-load experts.
-Next, we introduce Token Reroute to ensure a more balanced token-to-expert allocation.
-For tokens rejected by overflowed experts, we reselect underutilized experts based on the updated importance scores following Token Drop.
-As shown in Equation 12, masking the mappings to overflowed experts prevents these tokens from being reassigned to them, thereby encouraging their redistribution to other available experts.
-As illustrated in Algorithm 1, the process of masking and recomputing can be repeated iteratively, gradually achieving a more balanced token-to-expert assignment.
-Notably, the extra cost of Token Routing is minimal, as it occurs before tokens are assigned to each expert and operates solely on the lightweight importance matrix.
-Additionally, overflowed experts and tokens that have already selected $k$ experts are excluded from subsequent iterations, further minimizing the additional computational cost of the Top-$k$ operation.
+Next, we introduce Expanded Drop to ensure a more balanced token-to-expert allocation.
+A naive approach to rerouting under-selected tokens is to mask the mapping scores of overflowed experts and then reselect experts for these tokens.
+However, the reselection may still result in overflows, necessitating multiple rounds of selection and dropping, which increases latency.
+Moreover, the repeated selection and dropping substantially raise the cost of token-to-expert mapping.
+Expanded Drop adopts a simple yet effective strategy: for each token, it selects additional candidate experts.
+Given $m$ experts deployed on a single GPU, a token not only selects the top-$k$ experts based on gating scores, but also includes $m$ local experts (e.g., 8 experts per device under 8-way expert parallelism across 8 GPUs for a total of 64 experts) for substitution if the initially selected experts are overflowed.
+As a result, each token may select up to $m+k$ experts.
+The final selection is then refined as experts drop tokens as needed to satisfy capacity constraints.
+This makes no change in the token assignments in experts that are overflowed by the top-$k$ experts.
+Meanwhile, for under-utilized experts, the expanded top-$k+m$ candidate pool increases the likelihood of receiving additional tokens.
+After top-$k+m$ selection and dropping, there might exists tokens that select more than $k$ experts.
+Through empirical analysis (Appendix D), we choose not to enforce a constraint that limits each token to selecting at most $k$ experts, thereby removing the need to explicitly retain the top-$k$ experts at the end.
+Notably, the extra cost of token routing is minimal; the only difference lies in the negligible cost in the concatenation of the gating scores from either the top-$k$ or $m$ experts on the local device.
+Moreover, processing expanded tokens within local device eliminates inter-device communication.
 5 Experiments
 In this section, we conduct experiments under capacity-aware inference for MoE, with deployment details provided in Appendix A.
+5.1 Token Drop for High-load Experts
 Table 1: Performance comparison across different capacity factors and selection metrics (i.e., Order, Reverse Order, Random, and Score).
 The baseline operates without capacity constraints, represented as +$\infty$+ ∞.
-Due to inherent randomness, we report the average performance over multiple random seeds.
+We report the average performance over multiple random seeds.
 Method $\gamma$ OBQA PIQA RTE WinoGrande BoolQ ARC-C HellaSwag MMLU Avg.
 Baseline +$\infty$+ ∞ 45.6 80.1 53.7 71.2 74.7 54.5 79.4 52.5 64.0
-Order 3.0 43.6 75.8 53.1 71.2 74.4 50.5 78.6 51.7 62.4
-Reverse Order 44.8 76.1 50.9 71.0 74.8 52.8 78.7 51.7 62.6
-Random 44.0 76.3 52.0 71.3 75.0 53.5 78.9 51.8 62.9
-Score 45.2 79.9 51.6 71.7 74.4 55.2 79.2 51.9 63.6
 Order 2.0 42.0 71.5 53.1 71.2 74.2 49.5 76.6 48.4 60.8
@@
 Random 41.2 75.2 52.7 71.0 74.1 50.1 76.8 49.4 61.3
-Score 44.0 79.7 53.8 72.1 74.2 54.4 78.4 50.4 63.4
+Score 45.0 80.1 54.5 71.5 74.6 54.9 79.3 51.8 64.0
 Order 1.5 38.8 67.1 48.7 68.5 73.3 46.3 54.0 43.7 55.1
 Reverse Order 40.2 67.3 52.7 70.1 72.7 45.5 54.4 45.2 56.0
-Random 39.6 72.1 57.8 68.3 73.8 45.8 74.2 45.2 59.6
-Score 43.2 76.1 56.3 69.9 73.4 52.9 77.1 47.5 62.1
-Order 1.0 36.0 60.2 53.4 62.6 69.6 38.7 58.0 36.9 51.9
+Random 39.6 72.1 53.8 68.3 73.8 45.8 74.2 45.2 59.1
+Score 44.8 77.5 55.2 70.8 74.3 53.4 78.6 50.0 63.1
+Order 1.0 36.0 60.2 52.2 62.6 69.6 38.7 58.0 36.9 51.8
 Reverse Order 36.2 59.5 50.5 63.3 69.4 39.4 58.7 38.7 52.0
-Random 34.0 63.1 55.2 60.8 70.2 40.5 66.9 35.7 53.3
-Score 40.4 71.5 57.0 64.3 71.9 45.4 72.0 39.7 57.8
-5.1 Token Drop for High-load Experts
+Random 34.0 63.1 53.2 60.8 70.2 40.5 66.9 35.7 53.1
+Score 41.6 76.0 53.4 69.9 73.2 50.4 77.1 47.8 61.1
 Investigation on Token Drop Metrics
@@
 Consequently, we adopt "Score" as the default metric.
-(a) Single MoE Layer
-(b) End-to-End Model
-Figure 4: Speedup achieved by capacity-aware inference compared to the baseline without capacity constraints.
+Figure 4: Speedup of a single MoE layer compared to the baseline without capacity constraints, achieved through two capacity-aware inference methods: Token Drop and Expanded Drop.
+Figure 5: End-to-End Model Speedup
+Figure 6: Breakdown Analysis on OLMoE
 Efficiency Gains from Capacity-Constrained Inference
 We next explore the efficiency improvements achieved by imposing expert capacity.
-Specifically, we employ distributed inference using eight A30 GPUs, utilizing an 8-way Data Parallelism (DP) and 8-way Expert Parallelism (EP) strategy through the Megatron-LM framework Shoeybi et al.
-(2019).
-In Mixtral-8$\times$7B-Instruct model, each GPU hosts a single expert, whereas, in models like OLMoE-Instruct, multiple experts must be deployed on a single GPU (e.g., eight experts per GPU) due to GPU resource constraints.
-As illustrated in Figure 4, imposing constraints on expert capacity considerably accelerates inference across the four tested MoE models, in comparison to the baseline model without capacity limitations.
-The enhanced efficiency of each MoE layer (Figure 4 (a)) contributes to faster end-to-end inference (Figure 4 (b)).
-For instance, when the capacity factor $\gamma$ is set to 1.5, this configuration achieves a 1.94$\times$ speedup for a single MoE module and a 1.37$\times$ speedup for the entire end-to-end model.
-Furthermore, it is evident that the efficiency improves as the capacity decreases.
-Notably, the degree of acceleration is influenced by the numerical relationship between the number of experts and GPUs in Expert Parallelism.
-In Mixtral-8×8$\times$8 ×7B-Instruct, one-to-one deployment maximizes the effectiveness of capacity-aware inference.
-Conversely, in models where multiple experts share a single GPU, the acceleration gains are relatively modest.
-This occurs because the aggregated load from multiple experts diminishes the proportion of reduced load, which is achieved by limiting the straggler expert.
-Additionally, we believe that further system optimizations or allocating more GPUs for expert distribution would further enhance the potential of capacity-aware inference.
+Specifically, we employ distributed inference using eight H20 GPUs, utilizing an 8-way Data Parallelism (DP) and 8-way Expert Parallelism (EP) strategy through the Megatron-LM framework [36].
+Notably, in Mixtral-8$\times$7B-Instruct model, each GPU hosts a single expert, whereas, in models like OLMoE-Instruct, multiple experts must be deployed on a single GPU (e.g., eight experts per GPU) due to GPU resource constraints.
+As illustrated in Figure 4, imposing constraints on expert capacity through Token Drop and Expanded Drop, considerably accelerates inference across the four tested MoE models, in comparison to the baseline model without capacity limitations.
+The enhanced efficiency of each MoE layer (Figure 4) contributes to faster end-to-end model inference (Figure 6).
+Moreover, as the capacity factor $\gamma$ decreases, capacity-aware inference methods achieve significantly greater acceleration.
+Notably, the efficacy of acceleration is influenced by the numerical relationship between the total experts and the engaged GPUs in Expert Parallelism.
+As illustrated in Figure 6, for Mixtral-8×8$\times$8 ×7B-Instruct, deploying a single expert per GPU maximizes the effectiveness of capacity-aware inference.
+In this configuration, Token Drop and Expanded Drop achieve end-to-end model speedups of 1.87 $\times$ and 1.85$\times$, respectively, with $\gamma=1.5$.
+Conversely, deploying a greater number of experts on a single GPU results in more modest acceleration gains, as evidenced by the “8E/GPU” (OLMoE-Instruct and Deepseek-V2-Lite) the “10E/GPU” (Qwen1.5-MoE-Chat) in Figure 4 and Figure 6.
+This is because the aggregated load from multiple experts diminishes the proportion of reduced load, which is achieved by limiting the straggler expert.
+Therefore, it is anticipated that allocating more GPUs for expert distribution, thereby reducing the number of experts per GPU, would enhance the acceleration effect of capacity-aware inference.
+The breakdown analysis presented in Figure 6 demonstrates that our proposed capacity-aware inference methods substantially reduce the duration of expert computation, permutation and communication, while preserving a comparable cost for gate processing.
+Notably, the duration of permutation and communication increases when tokens are expanded across a range of global experts.
+This is due to the increased communication workload required to transmit expanded global tokens across various GPU devices.
+Consequently, these results underscore the necessity of restricting the expanded tokens to be processed by local experts.
 Figure 7: Analysis of dropped tokens with respect to capacity factors.
@@
 Given that expert capacity enforces MoE layers to discard overflowed tokens, we next establish the relationship between expert capacity and the corresponding number of dropped tokens.
-Specifically, for a capacity factor $\gamma$, the total proportion of dropped tokens is given by:
+For a capacity factor $\gamma$, the total proportion of dropped tokens is given by:
 $DT=\frac{\sum_{i=1}^{n}\text{ReLU}(N_{i}-\gamma\bar{N})}{\sum_{i=1}^{n}N_{i}},$ (12)
 where $\text{ReLU}(N_{i}-\gamma\bar{N})$ represent the number of dropped tokens for the $i$-th expert.
-Figure 5 visualizes the number of dropped tokens across different capacity factors for various test datasets, with a more detailed illustration provided in Appendix C.
-Although the most overloaded expert receives much more tokens than the expected number of tokens $\bar{N}$, regulating the maximum capacity has a limited impact on the overall number of accommodated tokens (e.g., reducing it by less than 20% under a capacity factor of 2.0), thereby maintaining competitive performance even after discarding overflow tokens.
-Moreover, dropping a small proportion of overflowed tokens can significantly reduce the latency caused by overloaded experts (e.g., dropping 12% of overloaded tokens promotes the inference speed by 94% in Mixtral-8$\times$7B-Instruct), highlighting the effectiveness of capacity-constrained inference in improving both performance and efficiency.
-Table 2: Comparison of Expert Drop, Token Drop and Token Reroute.
+Figure 7 visualizes the number of dropped tokens across different capacity factors for various test datasets, with a more detailed illustration provided in Appendix F.
+Although the most overloaded expert receives much more tokens than the expected number of tokens $\bar{N}$, regulating the maximum capacity has a limited impact on the overall number of accommodated tokens, thereby maintaining competitive performance even after discarding overflow tokens.
+Moreover, dropping a small proportion of overflowed tokens can significantly reduce the latency caused by overloaded experts (e.g., dropping 12% of overloaded tokens promotes the inference speed by 85% in Mixtral-8$\times$7B-Instruct), highlighting the efficacy of capacity-aware inference in improving both performance and efficiency.
+Table 2: Comparison of Expert Drop, Token Drop and Expanded Drop.
 The capacity factor $\gamma$ is set to 2.0 for OLMoE and DeepSeek-V2-Lite, and 1.5 for Qwen1.5-MoE-Chat and Mixtral-8×7B-Instruct.
-In Expert Drop, we only skip one out of eight experts for Mixtral-8$\times$7B-Instruct, and 10% of the lowest load experts for other models.
+For Expert Drop, each forward pass skips one out of eight experts for Mixtral-8$\times$7B-Instruct, and the bottom 10% of lowest load experts for other models.
 Model Method OBQA PIQA RTE WinoGrande BoolQ ARC-C HellaSwag MMLU GSM8K Avg.
@@
 Token Drop 47.8 77.9 64.6 69.2 80.0 57.2 79.7 51.5 32.4 62.3
-Token Reroute 47.2 79.4 66.3 70.5 80.9 57.1 80.3 52.3 34.4 63.2
+Expanded Drop 47.2 79.4 66.3 70.5 80.9 57.1 80.3 52.3 34.4 63.2
 Qwen1.5-MoE-Chat Baseline 42.4 79.9 72.9 70.0 81.3 54.1 80.4 59.8 52.0 65.9
@@
 Token Drop 40.4 78.8 72.6 69.1 80.9 53.0 80.0 59.3 51.9 65.1
-Token Reroute 43.4 79.1 72.6 69.6 81.1 53.4 80.3 59.3 52.1 65.6
+Expanded Drop 43.4 79.1 72.6 69.6 81.1 53.4 80.3 59.3 52.1 65.6
 DeepSeek-V2-Lite-Chat Baseline 45.4 81.4 72.6 75.5 82.9 61.0 81.5 57.3 66.4 69.3
@@
 Token Drop 45.2 78.3 72.6 74.0 83.2 59.3 80.9 57.3 62.7 68.2
-Token Reroute 45.4 79.4 73.3 75.4 83.2 60.4 81.5 57.2 64.1 68.9
+Expanded Drop 45.4 79.4 73.3 75.4 83.2 60.4 81.5 57.2 64.1 68.9
 Mixtral-8$\times$7B-Instruct Baseline 47.4 84.8 71.8 82.5 88.5 71.7 87.5 70.2 64.2 74.3
@@
 Token Drop 46.4 83.3 71.7 82.2 88.3 71.2 87.4 69.1 64.7 73.8
-Token Reroute 47.8 85.0 71.8 83.0 88.6 71.5 87.6 70.2 64.6 74.5
-5.2 Token Reroute to Low-load Experts
+Expanded Drop 47.8 85.0 71.8 83.0 88.6 71.5 87.6 70.2 64.6 74.5
+5.2 Expanded Drop to Low-load Experts
 Besides the experts overloaded with tokens, some low-load experts receive only a few tokens, raising important questions: Are these low-load experts redundant and removable, or should they be leveraged to balance token allocation?
-Recent works Lu et al.
-(2024); He et al.
-(2024) remove less important experts to improve efficiency, while our proposed Token Reroute increases their utilization by redistributing tokens for a more balanced assignment.
-Next, we investigate the significance of low-load experts and validate the effectiveness of Token Reroute.
+Recent works [29, 19] remove less important experts to improve efficiency, while our proposed Expanded Drop increases their utilization by redistributing tokens for a more balanced assignment.
+Next, we investigate the significance of low-load experts and validate the effectiveness of Expanded Drop.
 The Critical Role of Low-Load Experts
@@
 These findings indicate that retaining low-load experts better preserves the performance of MoE models.
-Effectiveness of Token Reroute
-We examine the effectiveness of utilizing low-load experts by rerouting overflowed tokens to them (i.e., Token Reroute) instead of simply discarding these tokens to meet the target capacity.
-Comparing Token Reroute with Token Drop, redistributing excess tokens to low-load experts enhances performance, yielding a 0.9% improvement in the average performance of Qwen1.5-MoE-Chat.
+Figure 8: Normalized expert load after Token Drop and Expanded Drop, with the capacity ratio set to $\gamma=1.5$.
+Effectiveness of Expanded Drop
+We examine the effectiveness of utilizing low-load experts by Expanded Drop instead of simply discarding these tokens to meet the target capacity.
+Comparing Expanded Drop with Token Drop, redistributing excess tokens to low-load experts enhances performance, yielding a 0.9% improvement in the average performance of Qwen1.5-MoE-Chat.
 Furthermore, considering the performance degradation observed in Expert Drop, our findings highlight the crucial role of low-load experts in maintaining model effectiveness.
+Expanded Drop overselects experts for each token to expand the selection scope to prompt a more balanced token-expert assignments.
+As shown in Figure 8, increasing the overselection ratio $m$ allows tokens to consider more candidate experts after being dropped from overflowed ones, thereby improving low-load expert utilization and balancing the expert load.
+5.3 Extension to Multi-modal Mixture of Experts
+In addition to applying capacity-aware inference to MoE models for language tasks, we also explore its effectiveness in multi-modal MoE settings.
+Specifically, we evaluate the OLMoE based MolmoE [12], across multi-modal benchmarks, including MME [15], MMBench [28], and SEED-Bench [27].
+Table 3: Capacity-aware inference for Multi-modality MoE models under different routing strategies. “Percep.” and “Cognit.” denote Perception and Cognition, respectively. $\gamma$ is set to 1.0.
+Method Strategy Percep. Cognit.
+Baseline 1358.1 269.6
+Token Drop Uniform 1248.4 245.4
+Expanded Drop 1307.6 273.6
+Token Drop Text First 1114.2 214.4
+Expanded Drop 1163.6 241.3
+Token Drop Image First 1346.5 288.9
+Expanded Drop 1362.1 297.1
+Given that the input sequence contains tokens from multiple modalities, we first investigate different token dropping strategies.
+Specifically, we first treat all tokens equally and drops those with the lowest scores (“Uniform”).
+Beyond this, considering the redundancy often found in image tokens, we also experiment with a strategy that prioritizes dropping image tokens before selectively removing text tokens (“Image First”).
+For comparison, we also consider drop text tokens first (“Text First”).
+As shown in Table 3, on the MME benchmark, this image-first strategy yields improved performance, highlighting the benefit of prioritizing dropping image tokens for the load balance in multi-modal MoE models.
+Figure 9: Multi-modal token assignments across different experts.
+Figure 10: Comparison on MMBench across six multimodal capabilities in Appendix B.
+Given the redundancy of image tokens and their large proportion in multi-modal tasks, we further investigate more aggressive capacity factors for Token Drop and Expanded Drop using the “Image First” strategy.
+Figure 10 demonstrates the effectiveness of Capacity-Aware Inference under low capacity constraints (i.e., $\gamma=0.5$).
+This is largely due to the high redundancy in image tokens [4], which allows a higher dropping ratio without significantly affecting performance.
+Meanwhile, the dominance of image tokens enables the use of very low capacity factors without significantly affecting text token retention, as illustrated in Figure 10.
+Therefore, dropping image tokens at higher ratios leads to more balanced token assignments and substantially improved inference efficiency.
+5.4 Ablation Study
+Model-Specific Imbalanced Property
+We explore the imbalance property in various models, such as OLMoE, DeepSeek, Qwen and Mixtral, which differ in both architectures (e.g., depth and width) and training strategies (e.g., training from scratch [31, 10] vs.
+training after upcycling [22, 39]).
+On the one hand, our findings in Appendix E reveal different training strategies result in significantly varying levels of imbalance.
+Specifically, MoE models trained from scratch exhibit a much higher degree of imbalance.
+For instance, OLMoE and DeepSeek-V2-Lite experience peak expert-wise token allocations exceeding $5\bar{N}$, whereas Qwen1.5-MoE and Mixtral are upcycled from dense language models, maintain a more balanced distribution, with peak expert-wise allocations staying below $3\bar{N}$.
+This is because upcycling initializes all experts with identical parameters [24], reducing divergence and promoting balanced training in the early stages.
+On the other hand, despite the widespread use of auxiliary balance loss in MoE training, it does not guarantee balanced token assignments across experts, as token distribution still varies significantly during inference on test data.
+This necessitates integrating expert capacity into the inference process.
 (a) OLMoE
@@
 Figure 11: Performance change as capacity factors decrease from 3.0 to 0.0.
-5.3 Ablation Study
-Model-Specific Imbalanced Property
-We explore the imbalance property in various models, such as OLMoE, DeepSeek-V2-Lite and Qwen1.5-MoE, which differ in both architecture (e.g., depth and width) and training strategies (e.g., training from scratch Muennighoff et al.
-(2024); DeepSeek-AI et al.
-(2024a) vs.
-training after upcycling Jiang et al.
-(2024); Team (2024b)).
-On the one hand, our findings in Appendix B reveal different training strategies result in significantly varying levels of imbalance.
-Specifically, MoE models trained from scratch exhibit a much higher degree of imbalance.
-For instance, OLMoE and DeepSeek-V2-Lite experience peak expert-wise token allocations exceeding $5\bar{N}$, whereas Qwen1.5-MoE and Mixtral are upcycled from dense language models, maintain a more balanced distribution, with peak expert-wise allocations staying below $3\bar{N}$.
-This is because upcycling initializes experts with the same parameters Komatsuzaki et al.
-(2023), while experts are randomly initialized when training from scratch, leading to a greater imbalance in token assignments Lo et al.
-(2024).
-On the other hand, despite the widespread use of auxiliary balance loss in MoE training, it does not guarantee balanced token assignments across experts, as token distribution still varies significantly during inference on test data.
-This necessitates integrating expert capacity into the inference process.
 Capacity Factor
@@
 We exclude values exceeding 3.0, as their performance closely aligns with capacity-agnostic scenarios.
-By analyzing the performance changes when decreasing the capacity factor, we find that setting $\gamma$ to 2.0 is sufficient to maintain performance comparable to the original models.
-This means allowing each expert to handle utmost twice the average number of tokens maintains the performance of capacity-free inference.
-However, maintaining the performance becomes challenging under low capacity factors (e.g., 1.0), as high-load experts experience significant token drops.
-Figure 7: Ablation study on the number of iterations for Token Reroute.
-Rerouting Round
-The Token Reroute Algorithm iteratively discourages tokens from selecting overflowed experts while encouraging their rerouting to other available experts, promoting a more balanced token distribution.
-As iterations progress, lower-score token-to-expert mappings are increasingly considered, further refining the redistribution process.
-Figure 7 illustrates the results across different iterations.
-In the first iteration, only excess tokens are dropped, while token rerouting becomes more prominent in subsequent rounds.
-The first two rounds of token rerouting lead to significant performance improvements, after which performance begins to saturate.
-We attribute this saturation to the selection of additional low-score token-to-expert mappings, which have a smaller impact compared to the dominant token assignments.
+By analyzing the performance changes when decreasing the capacity factor, we find that setting $\gamma$ to 1.5 is sufficient to maintain performance comparable to the original models.
+However, maintaining performance becomes challenging under excessively low capacity factors, as high-load experts are forced to drop a significant number of tokens.
 6 Conclusion
-In this paper, we first identify the imbalanced token-to-expert assignment in Mixture of Experts (MoE) and introduce the Straggler Effect in MoE inference, where the highest-load expert becomes the efficiency bottleneck, dictating overall latency.
-To address this issue, we propose Capacity-Aware Token Drop, which alleviates excessive loads on individual experts, and Capacity-Aware Token Reroute, which enhances the utilization of underutilized experts.
+In this paper, we identify the issue of imbalanced token-to-expert assignment in Mixture of Experts (MoE) models and introduce the Straggler Effect during inference, where high-load experts become efficiency bottlenecks and dictate overall latency.
+To address this problem, we propose Capacity-Aware Token Drop, which mitigates expert overload by enforcing strict capacity constraints.
+Additionally, to better utilize underloaded experts, we present Capacity-Aware Expanded Drop, which allows tokens to select additional experts on the same device while still respecting capacity limits, thereby improving expert utilization.
 Our findings and proposed methods offer valuable insights and effective strategies for improving MoE inference efficiency.
-Limitations
-Despite the progress we have made, our work still has certain limitations.
-First, our study focuses on a subset of MoE models, including OLMoE, Qwen1.5-MoE, DeepSeek-V2-Lite, and Mixtral.
-A promising direction for future work is to extend our analysis and methods to other architectures to further validate their effectiveness across a broader range of MoE models.
-Second, our proposed Capacity-Aware Token Drop and Capacity-Aware Token Reroute strategies could be integrated into the training process, potentially enhancing both performance and efficiency.
 Appendix A Implementation Details
@@
 We mainly focus on lightweight MoE models (less than 20B parameter budget).
-We conduct experiments on OLMoE Muennighoff et al.
-(2024), Qwen1.5-MoE Team (2024b), DeepSeek-V2-Lite DeepSeek-AI et al.
-(2024a) and Mixtral Jiang et al.
-(2024), due to their competitive performance and widespread adoption.
+We conduct experiments on OLMoE [31], Qwen1.5-MoE [39], DeepSeek-V2-Lite [10] and Mixtral [22], due to their competitive performance and widespread adoption.
 Datasets
 To evaluate model performance, we report normalized zero-shot or few-shot accuracy on the LM-Harness benchmark.
-The number of shots for each task is detailed in Table 3, which includes multiple tasks: ARC-C Clark et al.
-(2018), BoolQ Clark et al.
-(2019), HellaSwag Zellers et al.
-(2019), MMLU Hendrycks et al.
-(2021), OBQA Mihaylov et al.
-(2018), PIQA Bisk et al.
-(2019), RTE Wang et al.
-(2019), WinoGrande ai2 (2019) and GSM8K Cobbe et al.
-(2021).
-The evaluation code is based on EleutherAI’s LM Harness framework Gao et al.
-(2023).
+The number of shots for each task is detailed in Table 4, which includes multiple tasks: ARC-C [7], BoolQ [6], HellaSwag [43], MMLU [20], OBQA [30], PIQA [2], RTE [40], WinoGrande [1] and GSM8K [8].
+The evaluation code is based on EleutherAI’s LM Harness framework [16].
 Table 4: Experimental settings for evaluation tasks. “Norm” refers to the normalization performed with respect to the length of the input.
@@
 ARC-C 25 Accuracy (Norm)
-Appendix B Layer-wise Expert Load
+Appendix B Multi-Modal Tasks
+In the scope of this paper, multimodal tasks refer to those involving both vision and language modalities.
+We evaluate model performance using three representative benchmarks: MME, MMBench, and SEED-Bench, each targeting different aspects of multimodal understanding and reasoning.
+MME benchmark evaluates vision-language models along two dimensions: perception, which tests visual grounding and recognition, and cognition, which assesses reasoning abilities such as counting and relational understanding.
+It provides a fine-grained analysis of multimodal understanding.
+MMBench is a comprehensive benchmark designed to assess the general multimodal understanding ability of vision-language models.
+It evaluates model performance across six core capabilities: Coarse Perception (CP), Fine-grained Perception—including single-instance (FP-S) and cross-instance (FP-C), Attribute Reasoning (AR), Logical Reasoning (LR), and Relational Reasoning (RR).
+By covering both perception and reasoning-oriented tasks, MMBench provides detailed insights into the strengths and limitations of VLMs across diverse multimodal scenarios.
+SEED-Bench is a large-scale benchmark for evaluating the generative comprehension of Multimodal Large Language Models (MLLMs) across both image and video modalities.
+It includes 19K human-annotated multiple-choice questions spanning 12 evaluation dimensions, enabling objective and efficient assessment without human or GPT intervention.
+SEED-Bench reveals model limitations and maintains a public leaderboard to support fair comparison and future research.
+Component Content Token Count
+Image [Uncaptioned image] 576
+Text Prompt Is this artwork titled virgin and child with sts catherine, cecilia, barbara, and ursula? Please answer yes or no. 31
+Total 607
+Table 5: An example multi-modal query in MME benchmark, showing the dominant proportion of image tokens compared to text tokens.
+As shown in Table 5, these tasks typically introduce a large number of image tokens.
+When faced with imbalanced token-to-expert assignments, dropping redundant image tokens significantly improves load balancing.
+Moreover, due to the high redundancy among image tokens, dropping a portion of them has minimal impact on model performance.
+As in MME and MMBench, the Image-First variants of Token Drop and Expanded Drop also exhibit consistent effectiveness on SEED-Bench [27], maintaining strong performance even under low capacity factors such as $\gamma=0.5$.
+In addition to the redundancy in image tokens, Figure 10 shows that text tokens constitute only a small portion of the total token assignments.
+This allows the regulation of $\gamma$ to retain almost all text tokens under the Image-First strategy.
+Method $\gamma$ Inst. Attr. Inst. ID Inst. Interact. Inst. Loc. Inst. Count Scene Spatial Text Reasoning Overall
+Baseline $\infty$ 74.2 71.4 58.8 62.8 57.0 73.5 49.6 72.6 76.4 68.7
+Token Drop $0.5$ 70.4 67.8 60.8 57.8 51.8 71.5 43.5 58.3 71.3 64.9
+Expanded Drop 71.2 67.3 57.7 58.8 53.6 71.0 45.2 64.3 71.9 65.5
+Token Drop $1.0$ 73.6 70.2 58.8 62.6 56.9 72.7 48.1 61.9 73.4 68.0
+Expanded Drop 73.8 70.7 59.8 63.5 56.7 73.1 49.6 70.2 74.6 68.4
+Token Drop $1.5$ 73.5 71.2 58.8 62.9 57.1 73.1 48.7 65.5 74.6 68.3
+Expanded Drop 73.7 71.0 61.9 64.7 57.4 73.3 49.3 71.4 76.1 68.7
+Table 6: Token Drop and Expanded Drop strategies for multi-modal MoE models evaluated on SEED-Bench.
+Abbreviations: Inst.
+Attr.
+= Instance Attributes; Inst.
+ID = Instance Identification; Inst.
+Interact.
+= Instance Interaction; Inst.
+Loc.
+= Instance Localization; Inst.
+Count = Instance Counting.
+Appendix C Pseudocode for Token Drop and Expanded Drop
+We present the detailed pseudo-code for the algorithm implementations, as shown in Algorithm 1 and Algorithm 2. Both of these two methods adopt the selection metrics of “Score.”
+Algorithm 1 Token Drop
+input_tokens, num_tokens, num_experts, $k$, $\gamma$
+[…]
+scores $\leftarrow\text{Softmax}(\text{logits},\text{dim}=-1)$
+topk_scores, topk_indices $\leftarrow\text{TopK}(\text{scores},\text{k}=k,\text{dim}=1)$
+[…]
+[…]
+$\triangleright$ Drop Token
+expert_capacity $\leftarrow\gamma\cdot\frac{\text{num\_tokens}\times k}{\text{num\_experts}}$
+_, capacity_indices $\leftarrow$ TopK(topk_masked_scores, k=expert_capacity, dim=0, sorted=False)
+[…]
+[…]
+[…]
+return final_scores, final_map
+Algorithm 2 Expanded Drop
+input_tokens, num_tokens, num_experts, $k$, $\gamma$, local_expert_id_list
+[…]
+scores $\leftarrow\text{Softmax}(\text{logits},\text{dim}=-1)$
+topk_scores, topk_indices $\leftarrow\text{TopK}(\text{scores},\text{k}=k,\text{dim}=1)$
+$\triangleright$ Expand
+local_indices $\leftarrow$ torch.tensor(local_expert_id_list).repeat(num_tokens, 1)
+[…]
+local_scores $\leftarrow$ scores[:, local_expert_id_list]
+[…]
+[…]
+[…]
+$\triangleright$ Drop Token
+expert_capacity $\leftarrow\gamma\cdot\frac{\text{num\_tokens}\times k}{\text{num\_experts}}$
+_, capacity_indices $\leftarrow$ TopK(expanded_masked_scores, k=expert_capacity, dim=0, sorted=False)
+[…]
+[…]
+[…]
+return final_scores, final_map
+Appendix D Maximum Expert Selection
+Expanded Drop employs overselection and dropping to not only ensure load balance but also improve the utilization of underloaded experts.
+Although this process may allow some tokens to select more than $k$ experts, Table 7 shows that enforcing a strict $k$-expert limit is unnecessary: allowing additional expert selections has the potential of enhancing representational capacity, while rigid constraints introduce redundant computations.
+Table 7: Ablation study on limiting the maximum number of $k$ selected experts. “w/max” and “w/o max” indicate runs with and without this constraint, respectively. $\gamma$ is set to 1.0.
+$m$ Method OBQA PIQA RTE WinoGrande BoolQ ARC-C HellaSwag MMLU Avg.
+$2k$ w/ max 42.4 75.8 53.2 68.6 72.7 50.3 77.1 47.6 61.0
+w/o max 42.2 75.9 53.4 69.7 73.1 50.3 77.0 47.8 61.2
+$3k$ w/ max 42.0 75.6 53.4 69.6 72.7 50.3 77.0 47.4 61.0
+w/o max 42.0 75.6 53.8 69.8 72.9 50.3 77.1 47.6 61.1
+Appendix E Layer-wise Expert Load
 To analyze imbalanced token assignments, we measure the expert load for each expert by tracking the peak expert load while running MoE models on various test datasets.
-Figure 8, 9, 10 and 11 present the full results for the normalized layer-wise expert load for OLMoE, DeepSeek-V2, Qwen1.5-MoE, and Mixtral-8$\times$7B-Instruct, respectively.
+Figure 12, 13, 14 and 15 present the full results for the normalized layer-wise expert load for OLMoE, DeepSeek-V2, Qwen1.5-MoE, and Mixtral-8$\times$7B-Instruct, respectively.
+Appendix F Calculation Dropped tokens
+Based on Equation 12, we calculate the total number of dropped tokens across experts in each layer under different capacity factors, as illustrated in Figures 16, 18, 17, and 19.
 Figure 12: Layer-wise expert load in OLMoE-Instruct.
@@
 Figure 19: Dropped tokens with respect to capacity factors in Mixtral-8$\times$7B-Instruct.
-Appendix C Dropped tokens Calculation
-Based on Equation 13, we calculate the total number of dropped tokens across experts in each layer under different capacity factors, as illustrated in Figures 12, 14, 13, and 15.